On an average the box contain 10 articles is likely to have 2 defectives in a consignment of 100 boxes, the total number of boxes expected to have exactly one defective are
1
Expert's answer
2021-08-08T17:36:57-0400

Binomial distribution

"P=\frac{2}{10}=0.2, n=10"

since average "np=2" for 10 articles.

The probability of a box having exactly one defectives is"P(x\n=1)"


"=\binom{10}{x}P^x(1-P)^{10-x}, x=1 \\
=\binom{10}{1}(\frac{2 }{10 })^1(\frac{8 }{10 })^{9} \\
=0.268435456"


Thus, in a consignment of 100 boxes, "0.268435456\u00d7100=26.8435456" which is approximately 27 boxes are expected to have exactly one defectives articles.